# Sliding DFT-based Signal Recovery for Modulo ADC with 1-bit Folding Information

### v2

@@ III-G Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of the signal length $N_{0}$, window length $N$, window roll-off $\alpha$ ($0\leq\alpha\leq 1$) and oversamping factor $\mathrm{OF}=\frac{1}{\rho}$. The number of segments depends on the ratio $\frac{N_{0}}{N}$, and the parameter $\alpha$ dictates how large is the overlap between consecutive segments. For large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Looking at the complexity of the residual recovery algorithm at each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\in\mathbb{C}^{|\mathcal{S}_{i}|\times K}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(K|\mathcal{S}_{\max}|^{2}\right)$, where $|\mathcal{S}_{\max}|=\max_{i}|\mathcal{S}_{i}|$. Since $K\approx(1-\rho)N$ and $|\mathcal{S}_{\max}|\leq K$ under appropriate choice of $\mathrm{OF}$ and $\lambda^{\prime}$, then the complexity of computing $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is in $\mathcal{O}\left((1-\rho)^{3}N^{3}\right)$. Considering all the segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}(1-\rho)^{3}N^{2}}{1-\alpha/2}\right)$. Hence, the computational complexity of our proposed algorithm is linear in the signal length $N_{0}$.
Figure 4: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length ($N$) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=4$.
Figure 5: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length ($N$) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=8$.
Figure 6: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length ($N$) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=12$.
@@
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed sliding DFT-based recovery scheme can be improved by choosing a small oversampling factor (high $\rho$) that satisfies the sufficient condition in Theorem 1. However, reducing $\mathrm{OF}$ negatively impacts the MSE as shown in equation (17). The speed of the proposed sliding DFT-based recovery scheme can also be improved by choosing a small window length $N$. However, a small window length also needs a higher $\alpha$ to compensate for the increased spectral leakage. Finally, we note that the recovery algorithm in [14] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, i.e., rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}((1-\rho)^{3}N_{0}^{3})$. Since $N$ is typically chosen to be much smaller than $N_{0}$, our proposed sliding DFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [14]. Suppose the signal to be processed has $N_{0}=100,000$ samples. Our proposed algorithm with sliding window length $N=256$ and roll-off parameter $\alpha=0.25$ can unfold the whole signal at roughly $7.48\times 10^{-6}$ times the total time needed by the recovery method in [14] to unfold the whole signal.
@@ IV Numerical Results
In this section, we validate the MSE performance guarantees for the proposed algorithm established in the previous section by comparing the theoretical MSE results with the simulated performance. Numerical experiments are also conducted to gain additional insight on how the algorithm parameters affect the performance guarantees.
@@ IV-A Analysis of the Sufficient Condition for $b$
[…]
@@
Figures 4-6 depict the empirical evaluation of $\log_{2}\left(1+0.75\tilde{M}\right)$ for $\mathrm{OF}=4$, $\mathrm{OF}=8$, and $\mathrm{OF}=12$, respectively. This term accounts for the extra bit resolution needed by the proposed algorithm (in addition to the 3 bits in the first term of equation (16)) to satisfy the sufficient condition. Based on these evaluations, it can be seen that using a higher oversampling factor generally decreases the second term of equation (16). One insight that can be drawn from this observation is that a reduction in the modulo ADC resolution can be compensated by oversampling. Figures 4-6 also show that longer segment length and frequent crossing of the modulo thresholds increase $M$.
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=1$. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the discrete set $\{-2,-1,0,+1,+2\}$, and $T$ is the time difference between two consecutive pulses. Intuitively, the signal $f(t)$ is constructed by generating 100,000 raised-cosine pulses with random amplitudes at every symbol period $T$. This signal is fed to a low-resolution 4-bit modulo ADC with sampling rate $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters. Note that it is not practical to unfold $f(t)$ using the DFT-based methods described in [10, 13, 14, 11, 36] due to its large $N_{0}$ and its lack of periodicity.
[…]
@@
Figure 7: Simulated and theoretical MSE vs. oversampling factor of the proposed sliding DFT-based unfolding algorithm for $N=64$, $\alpha=0.25$, and $K_{\mathrm{SL}}=4,8$. Superimposed in the figure are the simulated and theoretical MSE curves of a conventional ADC.
@@
Figure 7 depicts the simulated MSE (in decibels) of the modulo ADC with sliding DFT recovery and its the theoretical MSE (based on Theorem 1) under the aforementioned settings. The oversampling factor is swept from $\mathrm{OF}=4$ to $\mathrm{OF}=40$. The numerical results for modulo ADC demonstrate that the derived MSE performance guarantee is accurate since it coincides with the simulated MSE for all $\mathrm{OF}$ and $K_{\mathrm{SL}}$ settings considered. It can be observed that larger $K_{\mathrm{SL}}$ increases the MSE of the sliding DFT unfolding method. This can be attributed to the widening of the passband region of the digital LPF used in the last step of the unfolding method. Consequently, more quantization noise are able to pass through the digital LPF.
@@
The theoretical MSE curve for $K_{\mathrm{SL}}=0$ (i.e. spectral leakage neglected) is also superimposed in Figure 7. It is evident from the figure that the slope of the $K_{\mathrm{SL}}=0$ MSE curve is steeper at the high $\mathrm{OF}$ regime compared to those of $K_{\mathrm{SL}}=4$ and $K_{\mathrm{SL}}=8$ MSE curves. This validates the asymptotic MSE growth rates mentioned in Section III. Nonetheless, the modulo ADC which uses our sliding DFT recovery still outperforms the conventional ones even if there is spectral leakage. The gap between the MSE curves of the modulo ADC and conventional ADC widens as $\mathrm{OF}$ is increased.
@@ V Summary and Future Directions
In this work, we considered a modulo ADC system in which the output samples are associated with a 1-bit folding information signal $c[n]$. The unfolded samples are obtained using the sliding DFT-based recovery scheme described in Section II-C. The advantage of our proposed unfolding scheme over existing Fourier-based recovery algorithms for modulo sampling is the significant reduction in the length of the observation window. This reduction in the observation time has a positive impact in the algorithm’s computational complexity. We also provide sufficient conditions on the oversampling factor and ADC resolution to attain a certain MSE performance guarantee. When these sufficient conditions are met, the MSE performance of the modulo ADC which uses our proposed recovery method is better compared to that of conventional ADCs. Moreover, we also showed that spectral leakage affects the asymptotic MSE growth rate of our sliding DFT recovery method. These theoretical results are substantiated by the numerical experiments conducted in Section IV.
It is worth mentioning that while the proposed unfolding scheme relies on the availability of 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be done by performing robust sparse recovery of $\underline{\tilde{z}^{(i)}_{w}}[n]$ using a sliding DFT version of the LASSO-$B^{2}R^{2}$[13]. This idea, together with its performance guarantees, is considered a potential subject of future work. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution are not tight. Hence, another future direction is to tighten these sufficient conditions. The modulo sampling architecture considered in this work did not take into account non-ideal modulo nonlinearities such as hysterisis and folding transients. Thus, we also intend to apply the sliding DFT recovery to generalized modulo sampling models and analyze how modulo hysterisis and folding transients affects the algorithm’s performance. Lastly, it is also of interest to apply this proposed recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers, radars, etc.).
@@ Appendix B Proof of Proposition 2
[…]
@@
This implies that
$\mathrm{OF}\geq\frac{3}{1-\frac{K_{\mathrm{SL}}}{N}}$
must also hold. The $\mathrm{OF}$ requirements are satisfied simultaneously if equation (13) holds.

### v3

@@ III-G Computational Complexity
We now derive the computational complexity of the proposed algorithm in terms of the signal length $N_{0}$, window length $N$, window roll-off $\alpha$ ($0\leq\alpha\leq 1$) and oversampling factor $\mathrm{OF}=\frac{1}{\rho}$. The number of segments depends on the ratio $\frac{N_{0}}{N}$, and the parameter $\alpha$ dictates how large is the overlap between consecutive segments. For large $N_{0}$, the number of segments to be processed by the algorithm is in $\mathcal{O}\left(\frac{N_{0}}{N\left(1-\alpha/2\right)}\right)$. Looking at the complexity of the residual recovery algorithm at each frame, the calculation of $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}\in\mathbb{C}^{|\mathcal{S}_{i}|\times K}$ is the most computationally-expensive operation. More precisely, this calculation is in $\mathcal{O}\left(K|\mathcal{S}_{\max}|^{2}\right)$, where $|\mathcal{S}_{\max}|=\max_{i}|\mathcal{S}_{i}|$. Since $K\approx(1-\rho)N$ and $|\mathcal{S}_{\max}|\leq K$ under appropriate choice of $\mathrm{OF}$ and $\lambda^{\prime}$, then the complexity of computing $\mathbf{V}_{\mathcal{S}_{i}}^{\dagger}$ is in $\mathcal{O}\left((1-\rho)^{3}N^{3}\right)$. Considering all the segments to be processed, the overall computational complexity of the proposed recovery method is in $\mathcal{O}\left(\frac{N_{0}(1-\rho)^{3}N^{2}}{1-\alpha/2}\right)$. Hence, the computational complexity of our proposed algorithm is linear in the signal length $N_{0}$.
@@
Based on the derived computational complexity of the proposed algorithm, the speed of the proposed sliding DFT-based recovery scheme can be improved by choosing a small oversampling factor (high $\rho$) that satisfies the sufficient condition in Theorem 1. However, reducing $\mathrm{OF}$ negatively impacts the MSE as shown in equation (17). The speed of the proposed sliding DFT-based recovery scheme can also be improved by choosing a small window length $N$. However, a small window length also needs a higher $\alpha$ to compensate for the increased spectral leakage. Finally, we note that the recovery algorithm in [14, 48] is a special case of our proposed recovery algorithm with $N=N_{0}$ and $\alpha=0$, i.e., rectangular window with no overlap. Hence, its complexity is in $\mathcal{O}((1-\rho)^{3}N_{0}^{3})$. Since $N$ is typically chosen to be much smaller than $N_{0}$, our proposed sliding DFT-based recovery is $\mathcal{O}\left(\frac{N_{0}^{2}(1-\alpha/2)}{N^{2}}\right)$ faster than the recovery algorithm in [14]. Suppose the signal to be processed has $N_{0}=100,000$ samples. Our proposed algorithm with sliding window length $N=256$ and roll-off parameter $\alpha=0.25$ can unfold the whole signal at roughly $7.48\times 10^{-6}$ times the total time needed by the recovery method in [14, 48] to unfold the whole signal.
@@ IV Numerical Results
Figure 5: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=4$.
Figure 6: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=8$.
Figure 7: Empirical evaluation of $\log_{2}\left(1+0.75M\right)$ as a function of segment length (NN) and $|\mathcal{S}_{i}|$ for $\mathrm{OF}=12$.
In this section, we validate the MSE performance guarantees for the proposed algorithm established in the previous section by comparing the theoretical MSE results with the simulated performance. Numerical experiments are also conducted to gain additional insight on how the algorithm parameters affect the performance guarantees. We also compare the proposed sliding DFT recovery against the HoD-based recovery.
@@ IV-A Analysis of the Sufficient Condition for $b$
We first investigate how the sufficient condition on $b$ (i.e., equation (16)) is affected by algorithm and input signal parameters such as segment length $N$, oversampling factor $\mathrm{OF}$, and number of non-zero elements of the 1-bit folding information signal $c[n]$. To this end, we consider three $\mathrm{OF}$ parameter settings ($\mathrm{OF}=4,\;8,\;12$) and three segment lengths (N=64, 128, 256N=64,\;128,\;256). We also consider three values of $|\mathcal{S}_{i}|$ ($|\mathcal{S}_{i}|=\frac{N}{32},\frac{N}{16},\frac{N}{8}$). For each combination of these parameter settings, we generate 100,000 realizations of $\mathbf{V}_{\mathcal{S}_{i}}$ by selecting $|\mathcal{S}_{i}|$ out of $N$ columns of $\mathbf{V}$. We then estimate the second term of equation (16) by calculating
$\tilde{M}=\max_{i\in\{1,\;2,\cdots,\;100,000\}}\|\mathbf{V}^{\dagger}_{\mathcal{S}_{i}}\mathbf{V}_{\mathcal{S}_{i}^{c}}\|_{\infty},$ (21)
@@
Figures 5-7 depict the empirical evaluation $\log_{2}\left(1+0.75\tilde{M}\right) fo$r $\mathrm{OF}=4$, $\mathrm{OF}=8$, and $\mathrm{OF}=12$, respectively. This term accounts for the extra bit resolution needed by the proposed algorithm (in addition to the 3 bits in the first term of equation (16)) to satisfy the sufficient condition. Based on these evaluations, it can be seen that using a higher oversampling factor generally decreases the second term of equation (16). One insight that can be drawn from this observation is that a reduction in the modulo ADC resolution can be compensated by oversampling. Figures 5-7 also show that longer segment length and frequent crossing of the modulo thresholds increase $M$.
@@
where $p_{\mathrm{rc}}(t)$ is a raised-cosine pulse with roll-off parameter $\beta=1$. For the numerical simulation, we set the filter span of the raised-cosine pulse to 20. The peak amplitude of the raised-cosine pulse $p_{\mathrm{rc}}(t)$ is set to unity. The pulse amplitude $A_{m}$ is drawn uniformly from the interval [0.5,+1][-0.5,+1], and $T=1$ second is the time difference between two consecutive pulses. Intuitively, the signal $f(t)$ is constructed by generating 50,000 raised-cosine pulses with random amplitudes at every symbol period $T$. Consequently, the maximum frequency is $\frac{\omega_{m}}{2}=\frac{\pi(1+\beta)}{T}$. The interval is chosen so that $f(t)$ has a non-zero mean value. We want to show that the algorithm can work on signals with non-zero average value. This signal is fed to a low-resolution 4-bit modulo ADC with sampling rate $\frac{1}{T_{\mathrm{s}}}=\frac{\mathrm{OF}\omega_{m}}{2\pi}$ to produce the quantized modulo samples $f_{\lambda^{\prime},\mathrm{q}}[n]$. The dither sequence and modulo threshold are configured according to the modulo ADC parameters.
For the sliding DFT-based recovery algorithm, we set the roll-off parameter of the tapered cosine window to $\alpha=0.5$ and window length $N=64$. Note that the length of the tapered cosine window is also the segment length and the DFT size of the algorithm. To investigate the impact of spectral leakage, we consider two different spectral leakage values: $\delta_{\mathrm{SL}}=\frac{\pi}{32},\;\frac{\pi}{16}$ (resp. $K_{\mathrm{SL}}=4,8$). The spectral leakage in the discrete frequency interval $\left(\rho\pi+\delta_{\mathrm{SL}},2\pi-\rho\pi-\delta_{\mathrm{SL}}\right)$ is assumed to be negligible. We will show later that this assumption is indeed valid for this simulation setup since the theoretical and simulated MSE performances coincide. Equation (18) is used to specify the modulo threshold $\lambda^{\prime}$. The simulated performance is evaluated by taking the squared difference between the true samples $f[n]$ and the signal unfolding output $\hat{f}[n]$ and averaging the results over all $N_{0}$ samples.
@@
Figure 8: Simulated and theoretical MSE vs. oversampling factor of the proposed sliding DFT-based unfolding algorithm for $N=64$, $\alpha=0.5$, $b=4$ and $K_{\mathrm{SL}}=4,8$. Superimposed in the figure are the simulated and theoretical MSE curves of a conventional ADC
@@
Figure 8 depicts the simulated MSE (in decibels) of the modulo ADC with sliding DFT recovery and its the theoretical MSE (based on Theorem 1) under the aforementioned settings. The oversampling factor is swept from $\mathrm{OF}=4$ to $\mathrm{OF}=50$. The numerical results for modulo ADC demonstrate that the derived MSE performance guarantee is accurate since it coincides with the simulated MSE for all $\mathrm{OF}$ and $K_{\mathrm{SL}}$ settings considered. It can be observed that larger $\delta_{\mathrm{SL}}$ increases the MSE of the sliding DFT unfolding method. This can be attributed to two things: (a) the widening of the passband region of the digital LPF used in the last step of the unfolding method, and (b) the increase in the $\lambda^{\prime}$. Consequently, higher quantization noise power would be present at the digital LPF output.
@@
The theoretical MSE curve for $K_{\mathrm{SL}}=0$ (i.e. spectral leakage neglected) is also superimposed in Figure 8. It is evident from the figure that the slope of the $K_{\mathrm{SL}}=0$ MSE curve is steeper at the high $\mathrm{OF}$ regime compared to those of $K_{\mathrm{SL}}=4$ and $K_{\mathrm{SL}}=8$ MSE curves. This validates the asymptotic MSE growth rates mentioned in Section III. Nonetheless, the modulo ADC which uses our sliding DFT recovery still outperforms the conventional ones even if there is spectral leakage. The gap between the MSE curves of the modulo ADC and conventional ADC widens as $\mathrm{OF}$ is increased. At $\mathrm{OF}=40$, the difference between the MSE incurred by the conventional ADC and the MSE incurred by the proposed sliding DFT-based recovery with $K_{\mathrm{SL}}=8$ is roughly 23 dB.
IV-C Comparison between the Sliding DFT-based recovery and HoD-based recovery
We now compare the performance of the proposed sliding DFT recovery against an existing unfolding method for modulo sampling. Note that it is not practical to unfold $f(t)$ in (22) using the DFT-based methods described in [10, 13, 14, 11, 36, 48] due to its large $N_{0}$ and its lack of periodicity. Moreover, the thresholding-based recovery method in [28, Algorithm 1] is unsuitable for the given $f(t)$, as it requires a minimum separation between folding times to function correctly. The low-sampling-rate variant [28, Algorithm 2] is likewise impractical due to its $\mathcal{O}(N_{0}^{2})$ computational complexity. As a benchmark, we compare our proposed sliding DFT-based recovery method against the HoD modulo recovery approach [6, 7].
Figure 9: Simulated MSE vs. oversampling factor of the sliding DFT-based recovery and HoD-based recovery for $b=3,4,5$.
Figure 9 shows the simulated MSE performance of the proposed sliding DFT-based recovery method compared to the HoD-based recovery for different values of $\mathrm{OF}$ and $b$. The parameters for the sliding DFT-based method are the same as in Section IV-B, except with $\delta_{\mathrm{SL}}=\frac{\pi}{32}$. For both methods, the modulo threshold $\lambda^{\prime}$ is set according to (18) and the test signal $f(t)$ is given in (22). When $b=4$ or $b=5$, the proposed method successfully unfolds the modulo ADC output across all $\mathrm{OF}$ values considered. In contrast, the HoD-based recovery shows very high MSE for $4<\mathrm{OF}<15$ when $b=4$, and for $4<\mathrm{OF}<8$ when $b=5$. These results highlight the advantage of the proposed sliding DFT-based recovery over HoD-based recovery in the low-sampling-rate, low-resolution regime. The performance advantage of our proposed sliding DFT-based recovery over HoD could be attributed to the availability of the 1-bit folding information.
We also note that both methods fail to unfold the modulo ADC output when $b=3$, suggesting that $b>3$ may be a necessary condition for the successful operation of the sliding DFT-based recovery in its current form. Further research is needed to confirm this conjecture. Alternatively, employing other quantizer architectures in conjunction with modulo sampling may enable sliding DFT-based unfolding at resolutions below $b=4$. For instance, the one-bit unlimited sampling approach in [55] uses a 1-bit sigma–delta quantizer together with modulo sampling to implement a 1-bit modulo ADC.
@@ V Summary and Future Directions
In this work, we considered a modulo ADC system in which the output samples are associated with a 1-bit folding information signal $c[n]$. The unfolded samples are obtained using the sliding DFT-based recovery scheme described in Section II-C. The advantage of our proposed unfolding scheme over existing Fourier-based recovery algorithms for modulo sampling is the significant reduction in the length of the observation window. This reduction in the observation time has a positive impact in the algorithm’s computational complexity. We also provide sufficient conditions on the oversampling factor and ADC resolution to attain a certain MSE performance guarantee. When these sufficient conditions are met, the MSE performance of the modulo ADC which uses our proposed recovery method is better compared to that of conventional ADCs. Moreover, we also showed that spectral leakage affects the asymptotic MSE growth rate of our sliding DFT recovery method. These theoretical results are substantiated by the numerical experiments conducted in Section IV. Under the same modulo threshold setting and $b=4,5$, the proposed sliding DFT recovery method also outperformed the HoD-based recovery at low values of the oversampling factor.
It is worth mentioning that while the proposed unfolding scheme relies on the availability of 1-bit folding information signal $c[n]$, the proposed recovery can be modified to work without $c[n]$. This can be done by performing robust sparse recovery of $\underline{\tilde{z}^{(i)}_{w}}[n]$ using a sliding DFT version of the LASSO-$B^{2}R^{2}$ [13] or a sliding DFT version of the OMP-based unfolding in [48]. This idea, together with its performance guarantees, is considered a potential subject of future work. We also note that the sufficient conditions derived for the oversampling factor and ADC resolution are not tight. Hence, another future direction is to tighten these sufficient conditions. The modulo sampling architecture considered in this work did not take into account non-ideal modulo nonlinearities such as hysteresis and folding transients. Thus, we also intend to apply the sliding DFT recovery to generalized modulo sampling models and analyze how modulo hysteresis and folding transients affect the algorithm’s performance. Lastly, it is also of interest to apply this proposed recovery algorithm to specific low-resolution signal processing systems (e.g., communication receivers, radars, etc.).
@@ Appendix B Proof of Proposition 2
$\;\leq$ $\;\frac{N}{\mathrm{OF}}\left(1+\frac{\|f(t)\|_{\infty}}{\lambda^{\prime}}\right)$ (25)
for all $i$. The first line follows from multiplying the number of level crossings in a single level and the number of folding levels. The second line comes from the trivial lower bound of the floor function, i.e., $\lfloor x\rfloor\leq x$. Since the bound in equation (B) is independent of $i$, this bound is also an upper bound for $\underset{i}{\max}\;|\mathcal{S}_{i}|$. Using equation (B), the following inequality holds:
$\frac{N}{N-\frac{N}{\mathrm{OF}}\left(1+\frac{\|f(t)\|_{\infty}}{\lambda^{\prime}}\right)-K_{\mathrm{SL}}}\geq\frac{N}{N-\underset{i}{\max}\;|\mathcal{S}_{i}|-K_{\mathrm{SL}}}.$ (26)
@@
This implies that $\mathrm{OF}\geq\frac{3}{1-\frac{K_{\mathrm{SL}}}{N}}$ must also hold. The $\mathrm{OF}$ requirements are satisfied simultaneously if equation (13) holds.